A movie theater has the goal of selling 1,500 tickets for a new movie during its first week.The first two days,it sold 30% of its ticket goal.What is the average number of tickets the theater needs to sell in the next five days in order to meet its goals?
Answer by rfer(16322)

1500*0.30=450
1500-450=1050
1050/5=210 per day